OpenStudy (anonymous):

Solve for x ln (2x - 1) = 0

OpenStudy (angela210793):

ln (2x - 1) =ln1 2x-1=1 2x=2 x=1...

OpenStudy (anonymous):

That's a good way to think about it. Alternatively you can use the definition of the logarithm: \[ln (2x - 1) = 0 \iff e^0 = 2x -1 \implies 2x - 1 = 1\]

OpenStudy (angela210793):

:)